HIGH RESOLUTION INFRARED SPECTRA AND VIB-ROTATIONAL ANALYSIS OF THE ν_{3} AND ν_{4} REGIONS OF CHLORINE NITRATE IN THE TEMPERATURE RANGE 190 TO 297 K

J.-M. FLAUD, \underline{J. ORPHAL}, CNRS, Laboratoire de Photophysique Moléculaire, 91405 Orsay Cedex, France; W. J. LAFFERTY, Optical Technology Division, NIST Gaithersburg, MD 20899, USA; M. BIRK, G. WAGNER, Institute for Optoelectronics, DLR, 82234 Oberpfaffenhofen, Germany.

Chlorine nitrate plays an important role in upper atmospheric chemistry. It was first detected in the stratosphere by Rinsland et al.^{a} by observation of the ν_{4} band a-type Q-branch. There have been a number of laboratory infrared spectroscopic studies on the ν_{4} band since that ime.^{b} Among them is the work of Bell, Duxbury and Stuart^{c} who used a tunable diode laser to study and model the a-type Q-branches of the two chlorine isotopic species ^{35}Cl and ^{37}Cl as well as that of the ν_{4} + ν_{9} − ν_{9} hot band. At room temperature, even at the high resolution of a diode laser, the bands are only partially resolved due to the overlapping of lines of the two main isotopic species and of lines from several strong hot bands. A beautifully resolved spectrum of the ν_{4} region over a limited range of J and K_{a} was obtained later by Xu, Blake and Sharpe^{d} from diode laser spectra of a molecular jet at about 7 K. In this work, 29 spectra of the region 750 to 900 cm^{−1} at temperatures ranging from 190 to 297 K and air pressures ranging from zero to 156 hPa were recorded at DLR with resolutions of 0.00094 to 0.0083 cm^{−1}. Using the ground-state constants of Müller et al.^{e} a vib-rotational analysis of the ν_{4} and ν_{3} fundamental bands as well the hot band, ν_{4} + ν_{9} − ν_{9}, of the most abundant ^{35}Cl isotopomer was carried out using the spectrum recorded at 191 K with the highest resolution and a zero air pressure. The corresponding cold bands for the lesser abundant isotope were partially assigned. Upper state rotational Hamiltonian constants were determined allowing precise modeling of the contours of the ν_{4} fundamental Q-branches over the temperature and pressure ranges studied as shown by comparisons with the other spectra.

aC. P. Rinsland et al., J. Geophys. Res. D 90, 7931 (1985).

bA. Goldman, C. P. Rinsland, J.-M. Flaud, and J. Orphal, J. Quant. Spectrosc. Rad. Transf. 60, 875 (1998).

cW. Bell, G. Duxbury, and D. D. Stuart, J. Mol. Spectrosc. 152, 283 (1992).

dS. Xu, T. A. Blake, and S. W. Sharpe, J. Mol. Spectrosc. 183, 228 (1996).

eH. S. P. Müller, P. Helminger, and S. H. Young, J. Mol. Spectrosc. 181, 363 (1997).